OpenStudy (agentjamesbond007):

Use U-Substitution ∫ 4/[(1+2x)^3]dx

sam (.sam.):

u= 1+2x du=2dx dx=du/2 Substitute back to equation, ∫ 4/u^3 (du/2) ∫ 2/u^3 du - 1/u^2 + c Substitute u=1+2x into equation, - 1/(1+2x)^2 + c

OpenStudy (anonymous):

∫ 4/[(1+2x)^3]dx = 4 ∫ dx /[(1+2x)^3] Let u = 1 + 2x -> du/2 = dx = 4/2 ∫ du/ u^3 = 2 ( -1/ 2u^2 ) = - 1/u^2 = -1/( 1 +x)^2 + C

OpenStudy (anonymous):

I mean -1/( 1 + 2x)^2 + C